OpenStudy (anonymous):

8i/4-2i The answer is -4 + 8i/5 but i don't how to get to it....HELP

OpenStudy (phi):

when you divide by a complex number, think Complex conjugate multiply top and bottom by 4+2i

OpenStudy (anonymous):

i did that and i do not know what i am doing wrong

OpenStudy (anonymous):

32i+16i^2

OpenStudy (anonymous):

thats for

OpenStudy (anonymous):

to

OpenStudy (anonymous):

the top

OpenStudy (phi):

(4-2i)(4+2i) notice that looks like (a+b)(a-b) which equals (a^2 - b^2) or you could multiply it out. what do you get for the bottom?

OpenStudy (anonymous):

12

OpenStudy (phi):

(4 - 2i) (4+ 2i) the long way: 4*4 + 4*2i -2i*4 - 2i*2i 16 +8i -8i -4i^2 8i - 8i is 0 16 - 4*i^2 by i*i is -1 16 - 4 * -1 16+4 20 shorter way: 4^2 + 2^2 (the i*i will make the - turn into a positive) = 20

OpenStudy (anonymous):

oh thank you i get it

OpenStudy (phi):

btw, I would not multiply out things up top right away. It is simpler to cancel things \[ \frac{ 8 i (4+2i)}{20} \] divide top and bottom by 4 \[ \frac{ \cancel{8}2 i (4+2i)}{\cancel{20}5} \] now multiply it out \[ \frac{8i +4i^2}{5} = \frac{-4+8i}{5} \]

OpenStudy (anonymous):

okay thanks i will look for that on the next few problems : )